use standard algorithm to calculate 5-5/4 -1/2


1
Expert's answer
2021-05-17T18:40:48-0400
"5-\dfrac{5}{4}-\dfrac{1}{2}=\dfrac{5\times4}{4}-\dfrac{5}{4}-\dfrac{1\times2}{4}"

"=\dfrac{20-5-2}{4}=\dfrac{13}{4}"

"5-\dfrac{5}{4}-\dfrac{1}{2}=\dfrac{13}{4}"